OpenStudy (jtug6):

Calc 2 problem someone mind lending a hand?

OpenStudy (jtug6):

\[\int\limits (\cos x)[(\ln(\sin(x)]/(\sin x)\]

OpenStudy (xapproachesinfinity):

what did you think about so far looks a good substitution no?

OpenStudy (xapproachesinfinity):

is this what you have \[\int \frac{\cos x \ln(\sin x) }{\sin x }dx\]

OpenStudy (xapproachesinfinity):

\(u=\sin x ~~~~~~\Longrightarrow ~~~~ du=\cos x dx\)

OpenStudy (xapproachesinfinity):

so now we get this integral after sub \[\int \frac{\ln u}{u}~~~du\]

OpenStudy (xapproachesinfinity):

yet another sub for this new integral \(t=\ln u ~~~\Longrightarrow ~~~dt=\frac{1}{u}du\)

OpenStudy (xapproachesinfinity):

i leave you with that.... finish it up

OpenStudy (xapproachesinfinity):

what does the new integrand look like?

OpenStudy (xapproachesinfinity):

answer: \[\ln^2(\sqrt{\sin x })\] see how it was found

OpenStudy (xapproachesinfinity):

+C of course i forgot that

OpenStudy (xapproachesinfinity):

actually error lol answer is : \[\frac{\ln^2(\sin x)}{2}+c\]

OpenStudy (jtug6):

nice! thanks. i just needed to know which technique would be acceptable in solving this, forgot how to derive natural logs but i see now. thanks a ton!

OpenStudy (xapproachesinfinity):

U sub was nice with this :)

OpenStudy (xapproachesinfinity):

welcome

OpenStudy (jtug6):

yup. issue i find is determining which technique to use. i guess you just try to use them all and see which one seems to work the best or may be the only one that works?

OpenStudy (xapproachesinfinity):

try to solve this \[\int\sqrt{\frac{x}{1-x^3}}~~dx \]

OpenStudy (xapproachesinfinity):

well it is about seeing what can work. i looked at sin and saw that cos at the top is its derivative so i knew substitution would work great once you do this more and more you develop the sense of knowing what works and what not

OpenStudy (xapproachesinfinity):

Training is the best way to do this :) more and more problems

OpenStudy (jtug6):

uhhh i'd imagine we'd rewrite 1-x^3 using trig sub?

OpenStudy (anonymous):

i would go the whole hog and let \(u=\log(\sin(x))\) making \(du=\frac{\cos(x)}{\sin(x)}dx\)

OpenStudy (xapproachesinfinity):

but if you differentiate that you get -3x^2 , which is something that you don't have

OpenStudy (jtug6):

yeah thats why i was mentioning remembering derivative of nautral logs haha @satellite73

OpenStudy (xapproachesinfinity):

yeah true :) i just realized that lol

OpenStudy (xapproachesinfinity):

@satellite73 wild sense of eyeballing

OpenStudy (jtug6):

so 1-x^3 using trig sub or was that a no? was thinking we could say 1-x^2 * x then rewrite 1-x^2 using x = sin theta perhaps?